The mean weight of 800 college students is 56 kg and the standard deviation is 5 kg. Assuming that the weight is normally distributed, determine how many students weigh: a. Between 60 kg and 70 kg. b. More than 68 kg.
1
Expert's answer
2021-10-26T12:47:31-0400

a.

"P(60<X<70)=P(X<70)-P(X\leq 60)"

"=P(Z<\dfrac{70-56}{5})-P(Z\leq \dfrac{60-56}{5})"

"=P(Z<2.8)-P(Z\leq 0.8)"

"\approx0.9974449-0.7881446\approx 0.209300"

"0.209300(800)=167"

167 students weigh between 60 kg and 70 kg.


b.


"P(X>68)=1-P(X\leq 68)"

"=1-P(Z\leq \dfrac{68-56}{5})=1-P(Z\leq 2.6)"

"\approx0.004661"

"0.004661(800)=4"

4 students weigh more than 68 kg.